Question: Payments of $670 are being made at the end of each month for 5 years at an interest rate of 8% compounded monthly. Calculate the Present value?
Expert's answer
1
Expert's answer
2019-11-21T10:50:04-0500

It is an ordinary annuity.

"PV=PMT\frac{1-(1+\frac{r}{n})^{-n*t}}{\frac{r}{n}}=670\frac{1-(1+\frac{0.08}{12})^{-12*5}}{\frac{0.08}{12}}=\$33043.35."